PROBING SOLVATION EFFECTS OF IBr^{−}(CO_{2})_{n} BY PHOTOELECTRON SPECTROSCOPY

\underline{ELISA MILLER}, LEONID SHEPS, RYAN CALVI, and CARL LINEBERGER, JILA, Department of Chemistry, University of Colorado at Boulder, Boulder, Co 80309.

We report photoelectron spectra of IBr^{−}(CO_{2})_{n} to understand the changes in IBr^{−} structure due to clustering. These experiments utilize anion photoelectron spectroscopy combined with velocity map imaging (VMI) to obtain photoelectron energy and angular distributions. Photodetachment of IBr^{−}(X^{2}Σ^{+}) with 300 nm radiation accesses three neutral states of IBr: the ground state (X^{1}Σ^{+}) and the next two excited states (^{3}Π_{2} and ^{3}Π_{1}). Known excitation energies between the neutral states allow a direct calculation of the EA of IBr, 2.53 eV. Observed vibrational progressions resulting from transitions to the excited 3Π states of IBr lead to equilibrium distances. Current experiments on IBr^{−}(CO_{2})_{n} will be compared to that of the bare anion to understand changes in the structure of the cluster with additional CO_{2} molecules.